What is the probability of finding a n = 2, â„“ = 1 electron between a0 and 2a0?
1
Expert's answer
2020-06-18T10:59:40-0400

The radial wave function of hydrogen atom is "R_{21}=\frac{1}{\sqrt{24a_0^3}}(\frac{r}{a_0})e^{-\frac{r}{2a_0}}"

The probability of finding a n = 2, â„“ = 1 electron between "a_0" and "2a_0" is

"P=\intop(R_{21})^2r^2dr =\intop(\frac{1}{\sqrt{24a_0^3}}(\frac{r}{a_0})e^{-\frac{r}{2a_0}})^2r^2dr"

"=\frac{1}{{24a_0^5}}\int r^4e^{-\frac{r}{a_0}}dr" between the limits "a_0" and "2a_0"

On solving the above integration, we get

P =0.04899